How do you solve this GDP question?

Coal Country Pittsburgh Steel Detroit Automotive Mining (CCM)(PS) $4,200 $0 (DA) $21,500 $9,000 Total Revenue $9,000 Intermediate $4,200 Goods Wages Interest Payments $1,000 $2,000 $3,700 $600 $300 $200 $10,000 $1,000 $500 $1,000 Rent $200 Profit $1,000 According to the table above, the Gross Domestic Product of this economy is Select one O a. $12,500 O b. $34,700 O c. $21,500 O d. $4,800
Answer #1

Answer

The correct answer is (c) $21500

According to value added Method , GDP = Gross Value added = Value Of Output - Intermediate consumption

Value of Output = Total revenue from all sectors.

Thus Here Value of Output = Total revenue from all sectors = 4200 + 9000 + 21500 = 34700

Intermediate consumption = Value of Intermediate goods across all sectors.

Hence, Intermediate consumption = Value of Intermediate goods across all sectors = 0 + 4200 + 9000 = 13200

Hence, GDP = Value Of Output - Intermediate consumption =  34700 - 13200

=> GDP = $21500

Hence, the correct answer is (c) $21500